mathslover (mathslover):

Are the angles \(\alpha\) and \(\beta\) given by, \(\large{\alpha = ( 2n + \frac{1}{2}) \pi \pm A}\) and \(\large{\beta = n \pi + (-1)^n (\frac{\pi}{2} - A)}\) same where \(\large{n \in I}\)

mathslover (mathslover):

There is ^ belongs to " n belongs to I " I means integers

mathslover (mathslover):

@ParthKohli @experimentX @amistre64 @phi @.Sam.

Parth (parthkohli):

\[n \in \mathbb Z \]

mathslover (mathslover):

Yeah ... sorry I didn't know the sign for "belongs to" in LaTeX . btw any hint for the question ?

mathslover (mathslover):

I don't think they are equal.

Parth (parthkohli):

I'm looking at it. What is \(A\)?

Parth (parthkohli):

is \(A\) any real number?

Parth (parthkohli):

Of course it is.

mathslover (mathslover):

anything, that is not given in the question ..

mathslover (mathslover):

I tried to take : firstly n as even i.e. n = 2k and the second time n = 2k + 1

mathslover (mathslover):

I solved that for beta for i) n = 2k and then ii) n = 2k + 1

Parth (parthkohli):

Yes, that's what I have done too!

mathslover (mathslover):

that gives the resultant one as: \[\large{\beta = (2k + \frac{1}{2})\pi \pm A}\]

Parth (parthkohli):

\[\large \beta = n\pi + \dfrac{\pi}{2} - A = \dfrac{2\pi n+ \pi}{2}{\pm A}{}\]Yesh

Parth (parthkohli):

\[\large \alpha = \dfrac{4\pi n + \pi}{2} \pm A\]BTW, that's - A in the above post

Parth (parthkohli):

Your argument seems valid.

Parth (parthkohli):

gotta go, exam tomorrow. :-(

mathslover (mathslover):

well it should be solved like this : \[\large{\textbf{For n = 2k , n is even}}\] \[\large{k(2\pi) + \frac{\pi}{2} - A}\] \[\large{\textbf{ For n = 2k + 1, n is odd }}\] \[\large{(2k) \pi + \frac{\pi}{2} + A }\] \[\large{\beta = (2k)\pi + \frac{\pi}{2} \pm A}\] \[\large{\implies n\pi + \frac{\pi}{2}\pm A = \beta}\]

mathslover (mathslover):

therefore Beta is not equal to alpha ... Best of luck @ParthKohli

mathslover (mathslover):

@ash2326